1. A basketball player who is 2.00 m tall is standing on the floor 10.0 m from the basket, as shown. If he shoots the ball at a 40.0 angle with the horizontal, at what initial speed must he throw so that it goes through the hoop without striking the backboard? Find the total time the ball is in flight. With what velocity (magnitude, direction) does the ball hit the hoop? The basket height is 3.05 m. State givens, etc. (As detailed as possible, please! Checking to see if doing arithmetic correctly and using the equations properly.) My answers Initial speed = 7.06 m/s. flight time = 2s, and velocity= (10.7m/s , 60 degrees)
Answer #1

Motion along Horizontal

R = v cos x * t

v t = 10 / cos 40

v t = 13.05...... (I)

Motion along vertical

3.05 - 2 = v sin 40 * t - 0.5 * 9.8* t^2

1.05 = 13.05 sin 40 - 4.9 t^2

Solving for t

t = 1.224 s

=======

Putting in (i)

v = 13.05 / 1.224

v = 10.663 m/s

========

Comment before rate in case any doubt, will reply for sure.. Goodluck